# DarkEQA: Benchmarking Vision-Language Models for Embodied Question Answering in Low-Light Indoor Environments

### v2

@@ I INTRODUCTION
Advances in vision-language models (VLMs) have significantly enhanced robotic capabilities, improving semantic scene understanding [peng2023openscene, liang2023open], spatial reasoning [dorbala2022clip, yang2025thinking], and vision-language-action (VLA) policies [zitkovich2023rt, driess2023palm, ahn2022can]. Numerous Embodied Question Answering (EQA) benchmarks have been proposed to assess this commonsense reasoning for embodied agents, largely assuming well-lit, ideal visual conditions [das2018embodied, majumdar2024openeqa]. However, household robots are often intended for 24/7 operation, which means they will frequently encounter low-light scenarios, such as nighttime, entering dark rooms or power blackouts. As robot deployment in varied environments grows, robust perception under these conditions is not an edge case but a core necessity [keller2020illumination]. Accordingly, benchmarks that explicitly stress-test embodied VLM reasoning under low illumination are essential to quantify real-world robustness. Nevertheless, acquiring large-scale, real-world low-light images with clean, paired annotations—ideally with corresponding well-lit reference views—is challenging and costly, which has hindered the construction of such benchmarks. As a result, existing benchmarks have largely overlooked systematic evaluation of VLM-based reasoning and perception under degraded illumination, limiting their ability to predict real-world robustness.
@@
To fill this evaluation void, we present DarkEQA, an open-source benchmark to systematically measure the perceptual primitives for embodied tasks under low-light conditions. The design of DarkEQA is primarily grounded in a physically based formulation, where all visual degradations are modeled at the RAW sensor data level (or in linear RGB space). This follows the physics of illumination and sensor noise to realistically simulate real-world Image Signal Processing (ISP) scenarios. Moreover, to ensure benchmark integrity and prevent potential data contamination [shumailov2024ai], all Question Answering (QA) pairs are deterministically generated via rule-based procedure, rather than depending on commodity VLM services. QA generation results in a family of queries targeting perceptual primitives, including from simple object recognition (e.g., “Is there a cushion in the image?”) to affordance reasoning (e.g., “I want to sleep, is this room suitable for this?”).
@@
DarkEQA provides 9.4k question–image pairs, a standardized evaluation protocol, and a public codebase to reproduce our low-light degradation pipeline. Our DarkEQA benchmarks a diverse set of vision–language models (VLMs), including both open- and closed-source systems [liu2024improved, li2024llava, wang2025internvl3, yang2025qwen3, hurst2024gpt]. We also evaluate a state-of-the-art low-light image enhancement (LLIE) model [feijoo2025darkir] as a preprocessing baseline. Our evaluation yields two observations. First, while humans can recognize structural scene information of input images from intensity contrast, all tested VLMs show a clear performance decline as the images degrade. Second, while LLIE preprocessing can improve performance under certain degradation levels, its effects are not consistently positive; in some cases, it yields limited gains or even leads to performance degradation, highlighting its practical limitations. Together, these results show that current VLM-based EQA pipelines remain brittle under low-light corruption, and that perceptual enhancement alone is insufficient as a general solution, motivating robustness-oriented evaluation and method development.
@@ II-A Embodied Question Answering Benchmark
Embodied Question Answering (EQA), first introduced by Das et al[das2018embodied], requires an agent to navigate and interact with an environment to answer a question. Early benchmarks primarily centered on static 3D scenes, such as ScanQA [azuma2022scanqa], to evaluate tasks like object identification and basic spatial relationships. OpenEQA [majumdar2024openeqa] is introduced to assess an agent’s exploration capabilities, posing diverse questions related to scene state, agent knowledge, and object attributes. Concurrently, a substantial body of research has focused on benchmarking deep spatial reasoning [zhu2025cospace, dang2025ecbench, zhang2025open3dvqa, song2025robospatial], evaluating complex object relationships [zhai2025memory]. Other works have pushed towards dynamic and procedural understanding, utilizing 3D scene graphs [armeni20193d, saxena2024GraphEQA, ali2025graphpad] or focusing on multimodal reasoning [hasegawa2025promqa, kamath2023gpr].
However, those existing EQA benchmarks often overlook real-world robustness. While NoisyEQA [wu2024noisyeqa] addresses query noise, robustness to adverse environmental conditions remains a significant gap. Notably, no current benchmark evaluates EQA in dark or low-light situations, which are common in the real world. We therefore introduce the first benchmark for indoor embodied question answering in dark environments to assess robustness under poor visibility.
@@ II-B Handling Low-Light Images
Recent research has explored two main directions for addressing the challenges of low-light visual perception. The first line of work targets robust recognition under low-light conditions, aiming to improve performance on specific vision tasks such as depth estimation, object detection, or pose estimation [gasperini_morbitzer2023md4all, wang2021regularizing, lee2023human, sasagawa2020yolo, spencer2020defeat]. Although these approaches demonstrate impressive robustness, they are typically constrained to single-task, highlighting a gap between low-light robustness in isolated perception and the embodied reasoning required in EQA. The second research stream focuses on low-light image enhancement (LLIE), where the goal is to improve the visual quality of dark images for human perception or downstream models [chen2018learning, lore2017llnet, lv2018mbllen, feijoo2025darkir, yuan2023clahe, li2015low]. These methods enhance brightness, contrast, and detail visibility using learning-based or physically inspired approaches. While LLIE methods improve visual quality, it remains unclear how they influence general embodied agents in low-light conditions. Therefore, we further explore whether LLIE can help EQA agents overcome the challenges they face in dark environments.
@@ III DarkEQA: Dataset Construction
Our DarkEQA is designed to evaluate VLMs’ recognition of core perceptual primitives from a single image-question pair under controlled low-light conditions. However, acquiring real-world low-light images with clean, paired annotations is challenging. To address this, we synthesize low-light images from the well-established indoor scene dataset (i.e., HM3D-Sem  [ramakrishnan2021hm3d]). This section describes the low-light image synthesis for benchmark inputs (Sec. III-A) and the EQA dataset construction process (Sec. III-B). A key feature of our work is a dataset construction pipeline designed for high reproducibility and expandability.
@@ III-A Low-Light Image Synthesis for Benchmark Inputs
Low-light images suffer from two distinct physical degradations. First, the reduced photon count leads to a fundamental loss of signal, which we term illumination degradation (i.e., exposure-value (EV) drop). Second, this weakened signal yields a low Signal-to-Noise Ratio (SNR), as sensor noise (e.g., shot, read, pattern, and quantization noise) becomes dominant relative to the remaining signal [wei2021physics]. To reproduce these conditions for benchmark inputs, we design a physics-based low-light synthesis pipeline. Specifically, across multiple degradation severities (L1–L5, increasing severity), we synthesize two paired low-light variants per original image: (i) A noise-free EV-drop variant and (ii) a physics-motivated variant with level-dependent sensor noise injection in the RAW domain, as in Fig. 3. This design enables disentangling the respective impacts of illumination degradation and sensor noise on perceptual peformance of VLMs.
@@ III-A1 Noise-free low-light image synthesis
First, we approximate linearization using gamma expansion. Let $x_{\mathrm{sRGB}}$ represent a sRGB pixel value in an input image and $x_{\mathrm{lin}}$ its linear form. Following [plotz2017benchmarking, brooks2019unprocessing], we compute
@@ III-A2 Physics-motivated low-light image synthesis
We synthesize realistic low-light images using a physics-based pipeline that combines ISP inversion/forward pass [brooks2019unprocessing] and raw-domain noise modeling [wei2021physics]. The process is shown in the upper branch of low-light image synthesis pipeline of Fig. 2.
@@
To obtain a camera-linear RAW image from $\mathbf{I}_{\mathrm{sRGB}}$, we invert the ISP following [brooks2019unprocessing]. We denote the unprocessing operator by ${u}(\cdot)$, and express the resulting Bayer RAW mosaic as
@@
Following the physics-based formation model of [wei2021physics], we inject four noise components into the camera-linear RAW signal. Let $\mathbf{B}$ denote the clean, mosaiced RAW image obtained from unprocessing. After converting $\mathbf{B}$ from normalized units to the sensor’s ADU domain, we sample a system gain $K$ log-uniformly from $[0.1,6.0]$. The noisy RAW image is then expressed as
@@
capturing the heavy-tailed distribution observed under extreme low-light [wei2021physics].
@@ III-B Dataset Construction
We build the dataset for evaluation upon a representative subset of 52 scenes from HM3D-Sem [ramakrishnan2021hm3d], selected for diversity and semantic richness. For each scene, we record a human-demonstrated navigation trajectory that systematically explores the environment to maximize spatial coverage. To generate the ground-truth QA pairs, we uniformly subsample the trajectory and select keyframes at a fixed time interval (e.g., one frame every 2,s), rendering their geometric and semantic modalities (e.g., RGB, depth, segmentation). We then use Algorithm 1 as deterministic procedure to automatically generate QA pairs from the pre-computed per-keyframe statistics. This approach ensures each question has a single, verifiable answer by filtering ambiguities (e.g., tiny objects), requires no manual annotation, and avoids potential data contamination by not using commodity VLM services. This entire process is fully reproducible.
@@ IV-B Baseline Models
We set the scenario of blind agents that produces an answer based on the question that requires visual information to answer [majumdar2024openeqa]. Even though our DarkEQA focuses on the VLM’s behavior according to illumination change and noise injection, we use the result of blind LLMs to catch the possible bias of our dataset while also testing how well the questions may be answered with an assumption of indoor environments. For the LLM choice, we report the results of GPT-4 [achiam2023gpt] and LLaMA-3.1-8B [dubey2024llama].
@@
We evaluate a range of VLMs across different parameter scales. For 7–8B models, we report results for LLaVA-1.6-7B [liu2024improved], LLaVA-OneVision-8B [li2024llava], InternVL3.5-8B [wang2025internvl3], and Qwen3-VL-8B [yang2025qwen3]. For larger-scale models ($\geq$ 30B), we additionally evaluate InternVL3.5-30B[wang2025internvl3] and Qwen3-VL-32B [yang2025qwen3] using the same respective series. Finally, we include GPT-4o [hurst2024gpt] as a upper bound.
@@
We use DarkIR [feijoo2025darkir] as our LLIE baseline throughout the evaluation for enhancing low-light images.
Figure 6: Summary of the evaluation results on our DarkEQA. Degradation level indicates the severity of low-light corruption: $L0$ corresponds to the original (well-lit) input, and higher levels ($L1\rightarrow L5$) denote progressively darker (lower-illumination) inputs. We evaluate a range of open-source VLMs (LLaVA [liu2024improved, li2024llava], InternVL [wang2025internvl3], and Qwen-VL [yang2025qwen3] series, 7B–32B). The shaded regions in (a) and (b) denote the minimum–maximum accuracy across models at each degradation level. (a) Impact of noise injection. (b) Impact of LLIE pre-processing. (c) Model-wise comparison. (d) Image samples enhanced with LLIE model. We include GPT-4 as a Blind-LLM baseline (evaluated without vision; gray dashed line) and GPT-4o [hurst2024gpt] as an upper-bound reference (black line).
@@ IV-C Results and Analysis
Given the observed performance degradation, we investigate whether pre-processing low-light images with a state-of-the-art Low-Light Image Enhancement (LLIE) model [feijoo2025darkir] can mitigate these issues. We apply LLIE models to the noise-added low-light images before feeding them into the VLMs. As illustrated in Fig. 6-(b), this approach yields mixed results. While we observe a significant accuracy improvement at more severe low-light levels (L4 and L5), performance decreases at moderate levels (L1–L3). This unstable behavior highlights the challenge of reliably enhancing low-light images across different levels of degradation. While current LLIE models enhance perceptual quality, the results suggest that current LLIE models may be biased to certain degradation levels as in Fig. 6-(d).

### v3

@@ I INTRODUCTION
Advances in vision-language models (VLMs) have significantly enhanced robotic capabilities, improving semantic scene understanding [27, 22], spatial reasoning [10, 40], and vision-language-action (VLA) policies [45, 11, 2]. Numerous Embodied Question Answering (EQA) benchmarks have been proposed to assess this commonsense reasoning for embodied agents, largely assuming well-lit, ideal visual conditions [9, 26]. However, household robots are often intended for 24/7 operation, which means they will frequently encounter low-light scenarios, such as nighttime, entering dark rooms or power blackouts. As robot deployment in varied environments grows, robust perception under these conditions is not an edge case but a core necessity [18]. Accordingly, benchmarks that explicitly stress-test embodied VLM reasoning under low illumination are essential to quantify real-world robustness. Nevertheless, acquiring large-scale, real-world low-light images with clean, paired annotations—ideally with corresponding well-lit reference views—is challenging and costly, which has hindered the construction of such benchmarks. As a result, existing benchmarks have largely overlooked systematic evaluation of VLM-based reasoning and perception under degraded illumination, limiting their ability to predict real-world robustness.
@@
To fill this evaluation void, we present DarkEQA, an open-source benchmark to systematically measure the perceptual primitives for embodied tasks under low-light conditions. The design of DarkEQA is primarily grounded in a physically based formulation, where all visual degradations are modeled at the RAW sensor data level (or in linear RGB space). This follows the physics of illumination and sensor noise to realistically simulate real-world Image Signal Processing (ISP) scenarios. Moreover, to ensure benchmark integrity and prevent potential data contamination [32], all Question Answering (QA) pairs are deterministically generated via rule-based procedure, rather than depending on commodity VLM services. QA generation results in a family of queries targeting perceptual primitives, including from simple object recognition (e.g., “Is there a cushion in the image?”) to affordance reasoning (e.g., “I want to sleep, is this room suitable for this?”).
@@
DarkEQA provides 9.4k question–image pairs, a standardized evaluation protocol, and a public codebase to reproduce our low-light degradation pipeline. Our DarkEQA benchmarks a diverse set of vision–language models (VLMs), including both open- and closed-source systems [23, 20, 36, 39, 16]. We also evaluate a state-of-the-art low-light image enhancement (LLIE) model [13] as a preprocessing baseline. Our evaluation yields two observations. First, while humans can recognize structural scene information of input images from intensity contrast, all tested VLMs show a clear performance decline as the images degrade. Second, while LLIE preprocessing can improve performance under certain degradation levels, its effects are not consistently positive; in some cases, it yields limited gains or even leads to performance degradation, highlighting its practical limitations. Together, these results show that current VLM-based EQA pipelines remain brittle under low-light corruption, and that perceptual enhancement alone is insufficient as a general solution, motivating robustness-oriented evaluation and method development.
@@ II-A Embodied Question Answering Benchmark
Embodied Question Answering (EQA), first introduced by Das et al[9], requires an agent to navigate and interact with an environment to answer a question. Early benchmarks primarily centered on static 3D scenes, such as ScanQA [5], to evaluate tasks like object identification and basic spatial relationships. OpenEQA [26] is introduced to assess an agent’s exploration capabilities, posing diverse questions related to scene state, agent knowledge, and object attributes. Concurrently, a substantial body of research has focused on benchmarking deep spatial reasoning [44, 8, 43, 33], evaluating complex object relationships [42]. Other works have pushed towards dynamic and procedural understanding, utilizing 3D scene graphs [4, 31, 3] or focusing on multimodal reasoning [15, 17].
However, those existing EQA benchmarks often overlook real-world robustness. While NoisyEQA [38] addresses query noise, robustness to adverse environmental conditions remains a significant gap. Notably, no current benchmark evaluates EQA in dark or low-light situations, which are common in the real world. We therefore introduce the first benchmark for indoor embodied question answering in dark environments to assess robustness under poor visibility.
@@ II-B Handling Low-Light Images
Recent research has explored two main directions for addressing the challenges of low-light visual perception. The first line of work targets robust recognition under low-light conditions, aiming to improve performance on specific vision tasks such as depth estimation, object detection, or pose estimation [14, 35, 19, 30, 34]. Although these approaches demonstrate impressive robustness, they are typically constrained to single-task, highlighting a gap between low-light robustness in isolated perception and the embodied reasoning required in EQA. The second research stream focuses on low-light image enhancement (LLIE), where the goal is to improve the visual quality of dark images for human perception or downstream models [7, 24, 25, 13, 41, 21]. These methods enhance brightness, contrast, and detail visibility using learning-based or physically inspired approaches. While LLIE methods improve visual quality, it remains unclear how they influence general embodied agents in low-light conditions. Therefore, we further explore whether LLIE can help EQA agents overcome the challenges they face in dark environments.
@@ III DarkEQA: Dataset Construction
Our DarkEQA is designed to evaluate VLMs’ recognition of core perceptual primitives from a single image-question pair under controlled low-light conditions. However, acquiring real-world low-light images with clean, paired annotations is challenging. To address this, we synthesize low-light images from the well-established indoor scene dataset (i.e., HM3D-Sem  [29]). This section describes the low-light image synthesis for benchmark inputs (Sec. III-A) and the EQA dataset construction process (Sec. III-B). A key feature of our work is a dataset construction pipeline designed for high reproducibility and expandability.
@@ III-A Low-Light Image Synthesis for Benchmark Inputs
Low-light images suffer from two distinct physical degradations. First, the reduced photon count leads to a fundamental loss of signal, which we term illumination degradation (i.e., exposure-value (EV) drop). Second, this weakened signal yields a low Signal-to-Noise Ratio (SNR), as sensor noise (e.g., shot, read, pattern, and quantization noise) becomes dominant relative to the remaining signal [37]. To reproduce these conditions for benchmark inputs, we design a physics-based low-light synthesis pipeline. Specifically, across multiple degradation severities (L1–L5, increasing severity), we synthesize two paired low-light variants per original image: (i) A noise-free EV-drop variant and (ii) a physics-motivated variant with level-dependent sensor noise injection in the RAW domain, as in Fig. 3. This design enables disentangling the respective impacts of illumination degradation and sensor noise on perceptual peformance of VLMs.
@@ III-A1 Noise-free low-light image synthesis
First, we approximate linearization using gamma expansion. Let $x_{\mathrm{sRGB}}$ represent a sRGB pixel value in an input image and $x_{\mathrm{lin}}$ its linear form. Following [28, 6], we compute
@@ III-A2 Physics-motivated low-light image synthesis
We synthesize realistic low-light images using a physics-based pipeline that combines ISP inversion/forward pass [6] and raw-domain noise modeling [37]. The process is shown in the upper branch of low-light image synthesis pipeline of Fig. 2.
@@
To obtain a camera-linear RAW image from $\mathbf{I}_{\mathrm{sRGB}}$, we invert the ISP following [6]. We denote the unprocessing operator by ${u}(\cdot)$, and express the resulting Bayer RAW mosaic as
@@
Following the physics-based formation model of [37], we inject four noise components into the camera-linear RAW signal. Let $\mathbf{B}$ denote the clean, mosaiced RAW image obtained from unprocessing. After converting $\mathbf{B}$ from normalized units to the sensor’s ADU domain, we sample a system gain $K$ log-uniformly from $[0.1,6.0]$. The noisy RAW image is then expressed as
@@
capturing the heavy-tailed distribution observed under extreme low-light [37].
@@ III-B Dataset Construction
We build the dataset for evaluation upon a representative subset of 52 scenes from HM3D-Sem [29], selected for diversity and semantic richness. For each scene, we record a human-demonstrated navigation trajectory that systematically explores the environment to maximize spatial coverage. To generate the ground-truth QA pairs, we uniformly subsample the trajectory and select keyframes at a fixed time interval (e.g., one frame every 2,s), rendering their geometric and semantic modalities (e.g., RGB, depth, segmentation). We then use Algorithm 1 as deterministic procedure to automatically generate QA pairs from the pre-computed per-keyframe statistics. This approach ensures each question has a single, verifiable answer by filtering ambiguities (e.g., tiny objects), requires no manual annotation, and avoids potential data contamination by not using commodity VLM services. This entire process is fully reproducible.
@@ IV-B Baseline Models
We set the scenario of blind agents that produces an answer based on the question that requires visual information to answer [26]. Even though our DarkEQA focuses on the VLM’s behavior according to illumination change and noise injection, we use the result of blind LLMs to catch the possible bias of our dataset while also testing how well the questions may be answered with an assumption of indoor environments. For the LLM choice, we report the results of GPT-4 [1] and LLaMA-3.1-8B [12].
@@
We evaluate a range of VLMs across different parameter scales. For 7–8B models, we report results for LLaVA-1.6-7B [23], LLaVA-OneVision-8B [20], InternVL3.5-8B [36], and Qwen3-VL-8B [39]. For larger-scale models ($\geq$ 30B), we additionally evaluate InternVL3.5-30B[36] and Qwen3-VL-32B [39] using the same respective series. Finally, we include GPT-4o [16] as a upper bound.
@@
We use DarkIR [13] as our LLIE baseline throughout the evaluation for enhancing low-light images.
Figure 6: Summary of the evaluation results on our DarkEQA. Degradation level indicates the severity of low-light corruption: $L0$ corresponds to the original (well-lit) input, and higher levels ($L1\rightarrow L5$) denote progressively darker (lower-illumination) inputs. We evaluate a range of open-source VLMs (LLaVA [23, 20], InternVL [36], and Qwen-VL [39] series, 7B–32B). The shaded regions in (a) and (b) denote the minimum–maximum accuracy across models at each degradation level. (a) Impact of noise injection. (b) Impact of LLIE pre-processing. (c) Model-wise comparison. (d) Image samples enhanced with LLIE model. We include GPT-4 as a Blind-LLM baseline (evaluated without vision; gray dashed line) and GPT-4o [16] as an upper-bound reference (black line).
@@ IV-C Results and Analysis
Given the observed performance degradation, we investigate whether pre-processing low-light images with a state-of-the-art Low-Light Image Enhancement (LLIE) model [13] can mitigate these issues. We apply LLIE models to the noise-added low-light images before feeding them into the VLMs. As illustrated in Fig. 6-(b), this approach yields mixed results. While we observe a significant accuracy improvement at more severe low-light levels (L4 and L5), performance decreases at moderate levels (L1–L3). This unstable behavior highlights the challenge of reliably enhancing low-light images across different levels of degradation. While current LLIE models enhance perceptual quality, the results suggest that current LLIE models may be biased to certain degradation levels as in Fig. 6-(d).